$SPAD/src/input dropt.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

\getchunk{license}
)clear all
--  Draw Options
draw(sin(1/x),x=-2*%pi..2*%pi,adaptive == true)
draw(sin(1/x),x=-2*%pi..2*%pi,adaptive == false)
draw(tan(x),x=-2*%pi..2*%pi,clip == true)
draw(tan(x),x=-2*%pi..2*%pi,clip == false)
draw(sin(x),x=-%pi..%pi,toScale == false)
draw(sin(x),x=-%pi..%pi,toScale == true)

--Other 2D Options:
draw(sec(x),x=-2*%pi..2*%pi,clip == [-2*%pi..2*%pi,-%pi..%pi])
draw(1/x,x=-10..10,clip == [-10..10])
draw(sin(x),x=-%pi..%pi,curveColor == 8.0)
draw(sin(x),x=-%pi..%pi,curveColor == bright red())
draw(sin(x),x=-%pi..%pi,pointColor == 3.0)
draw(sin(x),x=-%pi..%pi,pointColor == dim yellow())
draw(y**2 + y - (x**3 - x) = 0,x,y,range == [-2..2,-2..1])
p := ((x**2 + y**2 + 1) - 8*x)**2 - (8*(x**2 + y**2 + 1) - 4*x - 1)
draw(p = 0,x,y,range == [-1.0..11.0, -7.0..7.0])
seg1 : SEG FRAC INT := -3/2..3/2
range1 : LIST SEG FRAC INT := [seg1,seg1]
draw(x**2 + y**2 = 1,x,y,range == range1)
f1(t:DFLOAT):DFLOAT == 9*sin(3*t/4)
f2(t:DFLOAT):DFLOAT == 8*sin(t)
draw(curve(f1,f2),-4*%pi..4*%pi,unit == [2.0,0.5])
--Options used with 2D and 3D graphs:

g1(t:DFLOAT):DFLOAT == sin(5*t)
g2(t:DFLOAT):DFLOAT == t
draw(curve(g1,g2),0..2*%pi)
g1(t:DFLOAT):DFLOAT == sin(5*t)
g2(t:DFLOAT):DFLOAT == t
draw(curve(g1,g2),0..2*%pi,coordinates == polar)

m(u:DFLOAT,v:DFLOAT):DFLOAT == 1
draw(m,0..2*%pi,0..%pi,coordinates == spherical,title == "Sphere")
--3D Options:
colorFxn1(x:DFLOAT,y:DFLOAT):DFLOAT == x*sin(x)
draw(m,0..2*%pi,0..%pi,colorFunction == colorFxn1,title == _
  "color = x*sin(x)",coordinates == spherical)
colorFxn2(x:DFLOAT,y:DFLOAT):DFLOAT == x**2 - y**2
draw(m,0..2*%pi,0..%pi,colorFunction == colorFxn2,title == _
  "color = x**2 - y**2",coordinates == spherical)
colorFxn3(x:DFLOAT,y:DFLOAT,z:DFLOAT):DFLOAT == sin(x*z) + cos(y*z)
draw(m,0..2*%pi,0..%pi,colorFunction == colorFxn3,title == _
  "color = sin(x*z)+cos(y*z)",coordinates == spherical)
draw(curve(sin(t),cos(t),0),t=0..2*%pi,tubeRadius == .25)
draw(curve(sin(t),cos(t),0),t=0..2*%pi,tubeRadius == .25, tubePoints == 3)
draw(cos(x*y),x=-3..3,y=-3..3,var1Steps == 3, var2Steps == 3)
draw(cos(x*y),x=-3..3,y=-3..3,var1Steps == 9, var2Steps == 9)
draw(cos(x*y),x=-3..3,y=-3..3)

———-

References^{[1] nothing}

5